TTF2PK(1) TTF2PK(1) NAME ttf2pk−convertaTrueTypefontintoTeX’sPKformat SYNOPSIS ttf2pk[-q][-n]font-nameresolution ttf2pk-t[-q]font-name ttf2pk--version|--help DESCRIPTION ThisprogramconvertsaTrueTypefontintoaPKfile;thecreatedfontcanthenbeusedwithTEX orLATEX. AllTrueTypefontstobeusedmustberegisteredinaconfigurationfilecalledttfonts.map;it specifieshowtohandleeachfont.CJKV(Chinese/Japanese/Korean/oldVietnamese)subfontsas createdbyttf2tfmarealsosupported. ttf2pkalwaysassumes10ptasthedesignsizeforthewrittenTEXfont. PARAMETERS -qThisoptionalswitchmakesttf2pkquiet.Itsuppressesanyinformationaloutputexcept warninganderrormessages. -nUseonly‘.pk’astheextensioninsteadof‘.<resolution>pk’. -tTestfortheexistenceoffont-name.Returns0onsuccessandprintsoutthecorrespond-inglineinttfonts.map(providedthe-qswitchisn’tset). font-name TheTEXnameofthefont.ttf2pklooksthisnameupinaconfigurationfilecalled ttfonts.mapforfurtherinformationhowtoprocessthefont. resolution Theresolution,givenindotsperinch.Currentlythehorizontalresolutionisequaltothe verticalresolution.Thedesignsizeisalwaysassumedtobe10pt. --version Showsthecurrentversionofttf2pkandtheusedfilesearchlibrary(e.g.kpathsea). --helpShowsusageinformation. Environmentvariablesforfilesearchingaredescribedinthemanualpageofttf2tfm. THECONFIGURATIONFILE ttf2pkuses,similartodvips,afontdefinitionfilecalledttfonts.map.Theparametersspecified tottf2tfmareherepreserved—ttf2tfmwritesouttostandardoutput,asthelastline,aproper configurationentryforttfonts.map. Asanexample,acallto

ttf2tfmarial-pT1.enc-s0.25-P1-E0arials

willproducethefollowingline:

arialsarialSlant=0.25Pid=1Eid=0Encoding=T1.enc

Seettf2tfm(1)andafm2tfm(1)ofthedvipspackageforadetaileddescriptionofencodingfiles. Hereatablelistingthevariousttf2tfmparametersandthecorrespondingttfonts.mapentries: -s Slant -e Extend -p Encoding -f Fontindex -P Pid -E Eid -n PS=Yes -N PS=Only FreeTypeversion1.4 01-Mar-2000 1 TTF2PK(1) TTF2PK(1) -R Replacement -x Rotate=Yes -y Y-Offset Singlereplacementglyphnamesgiventottf2tfmwith‘-rold-glyphnamenew-glyphname’aredi-rectlyspecifiedwith‘old-glyphname=new-glyphname’.Theywillbeignoredifinsubfontmodeor ifnoencodingfileisgiven. Oneadditionalparameterinttfonts.mapisuniquetottf2pk:‘Hinting’,whichcantakethevalues‘On’or‘Off’.Somefonts(e.g.theCJKpartofcyberbit.ttf)arerenderedincorrectlyif hintingisactivated.Defaultis‘On’(youcanalsouse‘Yes’,‘No’,‘1’,and‘0’). Foradescriptionofsubfonts(i.e.,entriesoftheform‘<namestem>@<sfd-filename>@’)pleasere-fertottf2tfm(1). Theformatofttfonts.mapissimple.Eachlinedefinesafont;firstcomestheTEXfontname, thenitsTrueTypefontfilename,followedbytheparametersinanyorder.Caseissignificant (evenforparameternames);theparametersareseparatedfromitsvaluesbyanequalsign,with whitespacepossiblysurroundingit.ttf2pkreadsinttfonts.maplinebyline,continuinguntil theTEXfontspecifiedonthecommandlineisfound,otherwisetheprogramsexitswitherror code2.ThusyoucanuseanycharacterinvalidinaTEXfontnametostartacommentline. Youcanuse‘\’asthefinalcharacterofalinetoindicatethattheinputiscontinuedonthenext line.Thebackslashandthefollowingnewlinecharacterwillberemoved. ttf2pkwillabortifitcan’tfindorreadtheTEXfontmetricsfileofthegivenTEXfontname. RETURNVALUE Ifthecallwassuccessful,0willbereturned.Incaseoferror,thereturnvalueis1.Finally,ifthe fontcan’tbefoundinttfonts.map,2isreturned.Thissimplifiestheinclusionofttf2pkinto scriptslikemktexpkforautomaticfontgeneration. SEEALSO ttf2tfm(1),afm2tfm(1) AVAILABILITY ttf2pkispartoftheFreeTypepackage,ahighqualityTrueTyperenderinglibrary. AUTHORS WernerLEMBERG<wl@gnu.org>

Fre´ de´ ricLOYER<loyer@ensta.fr>

FreeTypeversion1.4 01-Mar-2000 2